A capillary tube of radius r can support a liquid of weight 6.28 × 10-4 N. if the surface tension of the liquid is 5 × 10-2 N/m. the radius of capillary must be

(a) 2.5 × 10-3 m (b) 2.0 × 10-4 m
(c) 1.5 × 10-3 m (d) 2.0 × 10-3 m

Grade:12

1 Answers

RANIVALLI
6 years ago
ANSWER : (D)
 
T=F/H
H=F/T
H=(6.28 x 10^-4)/(5 x 10^-2)
H=1.256 x 10^-2m
H=2T/Rdg
d=M/V
H=2TV/RMg
Mg=6.28 x 10^-4
V=(pi)(R^2)H
1=2T(pi)R/Mg
R=(6.28 x 10^-4)/(2 x 5 x 10^-2 x 3.14)
R=2.0 x 10^-3m.